### v2

# The Name of the Title Is Hope

Abstract.
Continual learning(CL) is a useful technique to acquire dynamic knowledge continually. Although powerful cloud platforms can fully exert the ability of CL,e.g., customized recommendation systems, similar personalized requirements for edge devices are almost disregarded. This phenomenon stems from the huge resource overhead involved in training neural networks and overcoming the forgetting problem of CL. This paper focuses on these scenarios and proposes a compact algorithm called LightCL. Different from other CL  methods bringing huge resource consumption to acquire generalizability among all tasks for delaying forgetting, LightCL  compress the resource consumption of already generalized components in neural networks and uses a few extra resources to improve memory in other parts. We first propose two new metrics of learning plasticity and memory stability to seek generalizability during CL. Based on the discovery that lower and middle layers have more generalizability and deeper layers are opposite, we Maintain Generalizability by freezing the lower and middle layers. Then, we Memorize Feature Patterns to stabilize the feature extracting patterns of previous tasks to improve generalizability in deeper layers. In the experimental comparison, LightCL  outperforms other SOTA methods in delaying forgetting and reduces at most 6.16$\times$ memory footprint, proving the excellent performance of LightCL  in efficiency. We also evaluate the efficiency of our method on an edge device, the Jetson Nano, which further proves our method’s practical effectiveness.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX11footnotetext: Equal contribution22footnotetext: Correspondence to chengfei@xidian.edu.cn
1. Introduction
The wave of deep learning has significantly boosted the deployment of the neural network from cloud platforms to edge devices(Lin et al., 2023, 2022). As the dynamic realistic environment requires the adaptive ability of models, a potential technique known as continual learning (CL) empowers these models to continually acquire, update, and accumulate knowledge on the device from the data by the training behavior. Although powerful cloud platforms can leverage CL  to adapt to service expansion, e.g., personalized recommendation systems(Lee et al., 2020), the similar and intense personalized requirements for edge devices are almost disregarded by academia and industry. This overlook stems from scarce resources in edge devices, such as limited computation, constrained storage, and insufficient memory, which significantly hinders the further application of CL. Hence, efficient CL  on edge devices becomes a considerable challenge.
Most existing CL  methods ignore edge scenarios and only concentrate on addressing the catastrophic forgetting problem(McCloskey and Cohen, 1989; Kirkpatrick et al., 2017) – model training on a new task tends to ”forget” the knowledge learned from previous tasks. Specifically, these methods explore two factors of generalizability(Wang et al., 2023), learning plasticity and memory stability, and make a trade-off between them to delay forgetting. Note that learning plasticity denotes the adaptation to the new task, while memory stability denotes the memory capacity of previous tasks. Hence, most of CL  methods(Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Kirkpatrick et al., 2017; Rusu et al., 2016) simultaneously train on the trace of historical information and new data for generalizability. It is well known that the training overhead of the neural network is already significant. Meanwhile, learning from history in CL  requires additional resources for the training process. Thus, the casual CL methods are unsuitable for resource-intensive scenarios. To mitigate this problem, compression techniques are inevitably introduced to the on-device CL. Especially, sparse learning(Han et al., 2015) has become a popular paradigm of the efficient CL(Wang et al., 2022; Ren and Honavar, [n. d.]). Although they(Wang et al., 2022; Ren and Honavar, [n. d.]) greatly reduce computation by the sparsity, intermediate feature maps cannot be effectively reduced, leading to an enormous memory footprint during training.
Let’s dive into the challenges of the on-device CL. Firstly, the forgetting problem is the first and foremost issue in CL, whether on edge devices or cloud platforms. Most CL methods train on historical information and new data together for generalizability to delay forgetting. This training behavior requires more significant resources than usual, including intermediate feature maps and gradients in memory, as well as additional computations for gradient calculations. This greatly hinders the efficiency. Secondly, the efficiency of CL heavily influences its prevalence in edge devices. The computation and memory footprint are two main factors that hinder the deployment. Meanwhile, memory rather than computation has become the primary bottleneck in AI applications, and the time taken for memory transfers greatly limits computational capacity(Gholami et al., 2024), which impedes the process of overcoming the forgetting during CL. Actually, the above two challenges are not orthogonal. The former inevitably interferes with the latter, and vice versa. This tangle prompts us to solve two challenges together. Considering the neural network has the potential to train in new tasks with the help of previous knowledge during CL, it absolutely has redundancy of generalizability. In other words, if we could evaluate the distribution of generalizability in the neural network during CL, we can compress the training redundancy with the over-strong generalizability and use just a few extra resources to maintain memory in less-generalized parts, which hopefully gives us a more favorable and efficient CL on edge devices.
Figure 1. Evaluation of accuracy and memory footprint on Split Tiny-ImageNet(Deng et al., 2009) under the TIL setting with sparsity ratio $90$%. The performance of LightCL outperforms the state-of-the-art efficient CL method SparCL(Wang et al., 2022) and reduces at most $6.16$$\times$ memory footprint compared with SparCLDER++.
Therefore, we start with the two factors of generalizability, learning plasticity and memory stability, and design their metrics to quantitatively assess the generalizability of neural networks during CL. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits significant variation. In particular, lower and middle layers have more generalizability, and deeper layers are the opposite. According to this conclusion, we Maintain Generalizability by easily freezing lower layers to delay forgetting and greatly reduce memory footprint and computation. Furthermore, for less-generalized parts in deeper layers, we regulate their feature maps, necessitating fewer resources, to Memorize Feature Patterns of previous tasks for the promotion of generalizability. The resource consumption of Memorize Feature Patterns is much smaller than its reduction of Maintain Generalizability. Thus, our method can achieve great performance with limited resources. Since our method achieves high efficiency only by exploring the network structure in CL, it can be easily combined with other compression techniques. As shown in Figure  1, the performance of LightCL outperforms the state-of-the-art efficient CL method SparCL and reduces at most $6.16$$\times$ memory footprint.
In summary, our work makes the following contributions:
  • It is the first study to propose two new metrics of learning plasticity and memory stability to quantitatively evaluate generalizability for efficient CL. We conclude that lower and middle layers have more generalizability, and deeper layers are the opposite.
  • Based on the evaluation, LightCL  efficiently overcomes catastrophic forgetting through Maintaining Generalizability and Memorize Feature Patterns.
  • Experiments illustrate that LightCL  outperforms state-of-the-art CL  methods in delaying the forgetting problem, and the memory footprint of LightCL  is reduced at most $6.16$$\times$.
The rest of the paper is organized as follows. Related work is reviewed in Section 2. In Section 3, we introduce the CL  setting related to this paper. We design the metrics and evaluate the generalizability in Section 4. According to the evaluation, we propose LightCL  in Section 5. Experimental analysis and conclusion are illustrated in Section 6 and Section 7, respectively.
2. RELATED WORK
Continual Learning Methods: CL  methods can be categorized into replay methods(Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Hayes et al., 2020; Cha et al., 2021), regularization-based methods(Kirkpatrick et al., 2017; Li and Hoiem, 2018; Zenke et al., 2017; Aljundi et al., 2018), and architecture-based methods(Mallya and Lazebnik, 2018; Rusu et al., 2016; Wang et al., 2020). Replay methods address catastrophic forgetting by replaying a subset of old data stored in storage. They must continually store samples when they meet new tasks, and bring extra computation and memory footprint for training on old data. Regularization-based methods are designed to restrict the updates of neural networks through regularization, preserving valuable information about previous tasks. Most of them(Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018) limit parameter changes necessary for old tasks, leading to extra consumption for recording the importance of every parameter. Architecture-based methods can prevent forgetting by distributing different model structures to each task. Similar work(Rusu et al., 2016) instantiates a new network for each new task, while others(Wang et al., 2020) divides a network into task-specific parts through pruning. Hence, most of them grow the model size continually or train on a whole dense model to learn new tasks, which causes significant resource overhead during the CL  process. Meanwhile, they must know the identity of the task before the model’s evaluation, which is unsuitable for edge scenarios.
Efficient Learning: Many efficient algorithms aim at the deployment of neural networks on edge devices, such as pruning(Han et al., 2016; Li et al., 2017), quantization(Gong et al., 2014) and etc. Pruning compresses networks by discarding weight redundancies, while quantization compresses the bit-width of weights. These methods mainly accelerate the inference procedure on edge(Shuvo et al., 2023), and degrade the ability of compressed models to accommodate dynamic environments. That’s because large training consumption hinders deployment development, such as massive usage of intermediate feature maps and gradients to memory footprint and more training FLOPs on computation. Although (Lin et al., 2022) can train in resource-limited settings, they focus more on training a single task and don’t consider catastrophic forgetting between multiple tasks. To cope with catastrophic forgetting of CL  in the edge device scenario, SparCL(Wang et al., 2022), combined with the dynamic pruning method, greatly reduces training FLOPs and delays the forgetting problem. However, SparCL still occupies massive memory footprints for intermediate feature maps.
3. CL Setting
In the CL  setting, tasks $\{1,2,...,T\}$ are executed sequentially during training, and data from previous tasks are inaccessible when training on the new task. Specifically, the t-th task represents $\mathcal{D}_{t}=\{(x_{t}^{i},y_{t}^{i})\}^{n_{t}}_{i=1}$, where $n_{t}$ is the sample number of t-th task and $(x_{t}^{i},y_{t}^{i})$ is input-label pair. $\theta_{i}^{t}$ represents the i-th layer of the neural network after training (or during) t-th task. The goal of CL  is to let the AI model perform well on all tasks when learning a new task. Overall performance is typically evaluated by average accuracy (AA)(Wang et al., 2023, 2022) on all tests after training. AA defined as $AA=\frac{1}{T}\sum_{i=1}^{T}a_{i}(\theta^{T})$, where $a_{i}(\theta)$ represents the testing accuracy of model $\theta$ on i-th task. There are two main CL  scenarios(Wang et al., 2022): 1) Task incremental learning (TIL), the identity $t$ of the task $\mathcal{D}_{t}$ are available during training and testing; 2) Class incremental learning (CIL), the task identity $t$ are only available during training. This paper evaluates CL  methods under TIL and CIL scenarios.
4. Analysis of Generalizability
In this section, we analyze the distribution of generalizability in the neural network during CL. Training consumption on generalized components in the neural network can be a kind of redundancy since they already have the ability to fit new tasks. That’s the motivation for evaluating generalizability. Considering generalizability means can be utilized for both previous and new tasks, we utilize memory stability (MS) and learning plasticity (LP) to represent these two different characters respectively. Similar work(Ramasesh et al., 2021) analyzes the forgetting parts in the neural network during CL. In fact, forgetting means strong learning plasticity and less memory stability, so generalizability and forgetting are not orthogonal and we cannot infer generalizability from it(Ramasesh et al., 2021). Thus, we need to design two metrics of LP and MS to quantitatively evaluate generalizability.
MS measures the loss of previous knowledge and LP measures adaptation to new knowledge. We evaluate them from the perspective of layers in the neural network. Specifically, we replace the parameters of i-th layer in the model $\theta^{T}$ with the parameters from the previous model $\theta^{T-1}$. The new model $\theta^{new}$ can be described as
(1) $\begin{gathered}\theta^{new}=\begin{cases}\theta^{new}_{j}=\theta^{T}_{j}&\text{if }j\neq i\\ \theta^{new}_{i}=\theta^{T-1}_{i}&\text{otherwise.}\end{cases}\end{gathered}$
Then, we separately test the accuracy of $\theta^{new}$ on (T-1)-th task and T-th task. Two metrics of LP and MS can be described as
(2) $\begin{gathered}\begin{cases}MS=a_{T-1}(\theta^{new})-a_{T-1}(\theta^{T})\\ LP=a_{T}(\theta^{new})-a_{T}(\theta^{T}).\end{cases}\end{gathered}$
If the value of MS is high, it indicates that the previous knowledge in i-th layer has been lost to a large extent, suggesting that the memory stability of i-th layer is weak. The low value of LP indicates that the previous i-th layer doesn’t have a good adaptation for the new tasks, meaning that the previous i-th layer has weak plasticity.
(a) Evaluation of MS
(b) Evaluation of LP
Figure 2. Evaluation of MS and LP on Split-TinyImageNet dataset with ResNet-18 (first row) and VGG16 (second row) under CIL setting. The X-axis is the index of each layer in the certain network. We operate two consecutive tasks for CL, containing the same number of classes, to evaluate the metrics of each layer. Different colors of lines represent different tests with different class numbers. All tests run $5$ times.
For more reasonable, we evaluate the above metrics using different class numbers for multiple tasks across $5$ runs on two typical backbones, ResNet-18(He et al., 2016) and VGG16(Simonyan and Zisserman, 2015). As shown in Figure  2, different numbers of classes exhibit a similar phenomenon during CL learning. Lower layers have small values of MS and high values of LP. In other words, the lower layers have high values of memory stability and learning plasticity, which exhibit good generalizability. The middle layers in ResNet-18 display good generalization ability, while the counterpart of VGG16 exhibits poor memory stability (second row in Figure 2(2(a))). It should be noted that LP suggests that middle layers already exhibit learning plasticity before being trained on the new task in Figure 2(2(b)), indicating that learning the new task will impair the stability of the middle layers and degrade the existing generalizability. This phenomenon proves the opinion that forgetting parts is not orthogonal to generalized parts. For the deeper layers, the values of MS are higher and the values of LP are lower. This suggests that the deeper layers are less-generalizability and forgetting parts. From the above analysis, we conclude that during CL, lower and middle layers have strong generalizability, and deeper layers have less generalizability. This conclusion guides the design of LightCL  in Section 5.
5. Methodology
LightCL, as shown in Figure  3, efficiently delays forgetting through Maintaining Generalizability and Memorize Feature Patterns. The whole pipeline of them is shown in Figure  LABEL:fig:pipeline_. All these methods are based on the discovery of generalizability in Section 4. We first introduce Maintaining Generalizability in Section 5.1 and then illustrate Memorize Feature Patterns in Section 5.2.
Figure 3. Illustration of each component in LightCL. Maintaining Generalizability and Memorize Feature Patterns work together to overcome catastrophic forgetting on resource-constrained edge conditions.
5.1. Maintaining Generalizability
According to the conclusion (Section 4) that lower and middle layers have strong generalizability, the training consumption on them can be a kind of redundancy. Hence, we easily freeze them to maintain generalizability without the training process, illustrated in the gray background section at the top of the Figure  3. Specifically, we must pretrain a network, freeze its lower and middle layers, and finally deploy it on the edge device for CL. Training the t-th task can be formulated as
(3) $\begin{array}[]{cl}\theta_{i}^{t}\leftarrow\theta_{i}^{t}+\alpha\nabla_{\theta_{i}^{t}}\mathcal{L}_{c}(f(\theta^{t},x_{t}),y_{t})\quad\text{s.t.}\quad i>FZ,\end{array}$
where FZ is the last layer index of frozen layers, $\mathcal{L}_{c}$ is the cross-entropy loss function, and function $f$ represents forward pass.
The frozen layers don’t involve backpropagation, so it does not need to store and compute many feature maps and gradients during training. We not only reduce memory footprint but also decrease training computation. As shown in Section 6.2.1, with the help of maintaining generalizability, LightCL  reduces at $6.16$$\times$ memory footprint and FLOPs are competitive beyond efficient CL  method SparCL. Meanwhile, in Section 6.3, it has greatly delayed forgetting in most CL  methods, which proves the universality of our method.
5.2. Memorize Feature Patterns
Based on the conclusion (Section 4) that deeper layers are less generalized and forgetting parts, we propose an efficient regulation method to memorize previous knowledge in deeper layers, shown in the pink background in the middle of the Figure  3. We go through the following three subsections to illustrate it.
5.2.1. How to Memorize Efficiently
Forgetting occurs because the gradient update of the current task does not consider generalizability among all tasks. Restricting the updating of vital weights of previous tasks, as the regularization-based method(Kirkpatrick et al., 2017) does, is the direct way to our goal. (Lin et al., 2022) points out that the deeper the layers, the more the memory footprint occupied by weights increases, while the size of feature maps becomes smaller and smaller. Our idea is that since both can reflect output patterns, we can memorize previous feature-extracting patterns by regulating important feature maps to overcome forgetting with a little extra consumption.
5.2.2. How to Select Reference
The intuitive idea is to regulate changes of vital feature maps produced by each sample of previous tasks during training for a new task. However, it is impossible to store samples continuously in resource-limited scenarios. As the number of tasks increases, so does the number of samples required. Alternatively, we regulate feature maps from the current sample instead of previous samples. That’s because regulating output patterns of feature maps is unrelated to the information of samples. The only effort is to recognize feature maps that require regulation. To realize this, we identify locations of feature maps vital to previous tasks by $\mathscr{l}_{1}$-norm(Li et al., 2017) with ratio $R$.
5.2.3. Detailed Process
After training the (t-1)-th task, neural network $\theta^{t-1}$ is generated. We get important positions of feature maps from the first batch of (t-1)-th task. Assume $(i,j)$ denotes j-th feature map in i-th layer to be position, all important positions are recorded in a set $I$. Before training new t-th task, we first select a few samples, defined as $\mathcal{M}$, from the new task to run it on the neural network $\theta^{t-1}$ and get feature map standard, called $FS$, to be referenced. The process is defined as $FS=F(\theta^{t-1},\mathcal{M},I)$, where $F$ is the function to get feature maps of important position $I$. Note that we only get $FS$ before training the new task without keeping acquiring during training, shown in the gray background part at the bottom of the Figure  3.
During training the t-th task, we integrate $M$ in each batch and get the corresponding feature maps. $FM$ is calculating from $FM=F(\theta^{t},\mathcal{M},I)$. The limiting changes of vital feature maps between $FS$ and $FM$ can be described as the regulation loss function
(4) $\begin{gathered}\mathcal{L}_{f}(FM,FS)=\beta\sum_{(i,j)\in I}(FM_{i,j}-FS_{i,j})^{2}.\end{gathered}$
The entire loss function $\mathcal{L}$ is the combination of cross-entropy loss function $\mathcal{L}_{c}$ and the regulation loss $\mathcal{L}_{f}$, described as follow
(5) $\begin{gathered}\mathcal{L}(\theta^{t})=\mathcal{L}_{c}(f(\theta^{t},x_{t}),y_{t})+\mathcal{L}_{f}(FM,FS).\end{gathered}$
The backward process of $\mathcal{L}$ is described as
(6) $\begin{array}[]{cl}\theta_{i,j}^{t}\leftarrow\theta_{i,j}^{t}+\alpha\nabla_{\theta_{i,j}^{t}}\mathcal{L}(f(\theta^{t}))\quad\text{s.t.}\quad i>FZ.\end{array}$
After the t-th task, we still use the first batch to get the important position $I^{\prime}$ of the current task with ratio $R$ and combine it with the last version, $I=I\lor I^{\prime}$. The whole process of the algorithm is shown in Algorithm 1. As shown in Section 6.3, Memorize Feature Patterns successfully delays forgetting with a few extra resources. The combination of Maintaining Generalizability and Memorize Feature Patterns outperforms other state-of-the-art   CL  methods(see in Table 1). Different from other replay methods, we are not required to store samples in storage, as they are only utilized during the current training. This also enhances privacy security.
Input :  Important Position Set $I$; Mem Sample $\mathcal{M}$; Function of Getting Feature Map $F$; Function of Neural Network $f$; Cross-entropy loss $\mathcal{L}_{c}$; Feature Regularization $\mathcal{L}_{f}$; Selecting ratio $R$
1 Freezing lower layers of a pre-trained neural network $\theta$
2 for $t=1,...,T$ do
3       $\mathcal{M}\leftarrow sample(\mathcal{D}_{t})$
4       $FS\leftarrow F(\theta,\mathcal{M},I)$
5      for iteration $Batch$ from $\mathcal{D}_{t}$ do
6             $(x,y)\leftarrow Batch$
7             $FM\leftarrow F(\theta,\mathcal{M},I)$
8             $\mathcal{L}\leftarrow\mathcal{L}_{c}(f(\theta,x),y)+\mathcal{L}_{f}(FM,FS)$
9             $\theta\leftarrow Update(\mathcal{L},\theta)$
10      $(x,y)\leftarrow$ last $Batch$ from $\mathcal{D}_{t}$
11       $FM\leftarrow F(\theta,x)$
12       Select Important Position $I^{\prime}$ with ratio $R$ from FM
13       $I=I^{\prime}\lor I$
14
Algorithm 1 LightCL  Algorithm
6. Experiment
Table 1. Comparison with CL  methods on pre-trained model. Classification results on ResNet-18 for standard CL  benchmarks, averaged across $5$ runs. LightCL  is outstanding beyond other CL  methods with fewer resources.
Method Buffer Sparsity Split CIFAR10 Split TinyImageNet
CIL()($\uparrow$)( ↑ ) TIL()($\uparrow$)( ↑ ) FLOPs Mem CIL()($\uparrow$)( ↑ ) TIL()($\uparrow$)( ↑ ) FLOPs Mem
size ×$\times 10^{15}(\downarrow)$ (MB)()($\downarrow$)( ↓ ) ×$\times 10^{16}(\downarrow$) (MB)()($\downarrow$)( ↓ )
JOINT / 0.0 94.67$\pm$0.15 99.04$\pm$0.07 37.5 153.4 61.38$\pm$0.12 83.14$\pm$0.52 120 357.9
SGD 34.21$\pm$3.06 86.19$\pm$1.32 7.5 153.4 10.63$\pm$0.36 34.24$\pm$0.84 12 357.9
EWC(Kirkpatrick et al., 2017) / 0.0 40.42$\pm$3.81 89.77$\pm$3.01 7.5 196.0 11.17$\pm$1.55 34.70$\pm$1.55 12 400.5
PNN(Rusu et al., 2016) / 94.62$\pm$0.52 10.5 323.9 / 73.94$\pm$3.56 26.4 528.4
FDR(Benjamin et al., 2019) 15 0.0 25.64$\pm$2.52 85.54$\pm$3.08 11 185.4 10.80$\pm$0.71 34.42$\pm$0.38 17.7 485.7
ER(Chaudhry et al., 2019) 54.44$\pm$2.82 92.25$\pm$1.92 11 185.4 11.82$\pm$0.31 36.76$\pm$0.57 17.7 485.7
DER++(Buzzega et al., 2020) 40.75$\pm$2.55 85.68$\pm$4.35 14.5 217.3 7.98$\pm$1.84 35.88$\pm$0.41 23.3 613.5
SparCLER(Wang et al., 2022) 15 0.9 36.98$\pm$2.42 85.33$\pm$2.49 1.1 106.6 10.01$\pm$0.55 33.01$\pm$0.78 1.8 407.0
SparCLDER++(Wang et al., 2022) 35.58$\pm$0.84 85.28$\pm$1.00 1.5 140.6 9.12$\pm$1.02 31.86$\pm$1.77 2.3 536.8
LightCL 15 0.0 55.53$\pm$0.35 96.28$\pm$1.08 5.4 92.8 12.79$\pm$0.30 37.39$\pm$0.63 8.6 128.0
0.9 38.31$\pm$3.93 86.88$\pm$1.69 2.3 25.0 8.95$\pm$0.35 33.94$\pm$0.52 3.7 87.1
‘/’ indicates the corresponding item is not needed (Buffer is not required in JOINT, SGD, EWC, PNN) or cannot run (PNN cannot run
under CIL settings). FLOPs and Mem are calculated theoretically with the precision to one decimal place.
Table 2. Comparison with CL  methods from scratch. LightCL  still outperforms other CL  methods.
Method Split CIFAR10 Split TinyImageNet
CIL()($\uparrow$)( ↑ ) TIL()($\uparrow$)( ↑ ) CIL()($\uparrow$)( ↑ ) TIL()($\uparrow$)( ↑ )
JOINT 92.20$\pm$0.15 98.31$\pm$0.12 59.99$\pm$0.19 82.04$\pm$0.10
SGD 21.00$\pm$1.61 65.78$\pm$3.52 7.28$\pm$0.04 24.99$\pm$0.27
EWC 21.32$\pm$0.17 66.13$\pm$0.17 8.58$\pm$0.22 24.47$\pm$0.30
PNN / 95.13$\pm$0.72 / 61.88$\pm$1.00
FDR 26.19$\pm$3.14 77.61$\pm$0.68 8.58 $\pm$0.21 25.20$\pm$0.29
ER 35.90$\pm$1.53 82.40$\pm$1.64 8.76$\pm$0.23 26.29$\pm$1.18
DER++ 36.25$\pm$1.99 82.05$\pm$1.63 6.53$\pm$1.44 27.21$\pm$1.85
SparCLER 30.09$\pm$1.60 75.12$\pm$0.63 7.70$\pm$0.13 25.24$\pm$0.55
SparCLDER++ 34.36$\pm$1.32 78.56$\pm$2.88 8.16$\pm$0.58 26.53$\pm$0.54
LightCL 37.86$\pm$2.16 85.69$\pm$0.98 9.03$\pm$0.16 33.74$\pm$0.68
LightCL0.9 35.72$\pm$2.67 83.37$\pm$1.10 8.76$\pm$0.63 36.86$\pm$0.29
6.1. Experiment Setting
Datasets: We evaluate LightCL  on Split CIFAR-10(Krizhevsky et al., 2009) and Split Tiny-ImageNet(Deng et al., 2009). To compare with other CL  methods, we strictly follow (Buzzega et al., 2020; Wang et al., 2022) by splitting CIFAR-10 and Tiny-ImageNet into $5$ and $10$ tasks, and each task including $2$ and $20$ classes, respectively. For Split CIFAR-10 and Tiny-ImageNet, we train each task for $50$ and $100$ epochs with a batch size of $32$.
Baseline Mehtods: We select typical CL  methods from three categories, mentioned in section 2, including regularization-based (EWC(Kirkpatrick et al., 2017)), replay (DER(Buzzega et al., 2020), ER(Chaudhry et al., 2019), FDR(Benjamin et al., 2019)), and architecture-based (PNN(Rusu et al., 2016)) methods. Since these above methods do not consider the edge scenario, we also compare our method against the state-of-the-art   efficient CL  method SparCL(Wang et al., 2022) combined with the replay method DER++ and ER. We use SGD without any CL  method as a lower bound and train all tasks jointly (JOINT) as an upper bound.
Metrics: We use AA, mentioned in 3, to evaluate the performance of overcoming forgetting. Moreover, we follow (Wang et al., 2022) to theoretically assess the training FLOPs and peak memory footprint (including feature maps, gradients, and model parameters during training) for the illustration of resource usage.
Setting of LightCL: We use a pre-trained ResNet-18 model (pre-trained on the ImageNet32323232$\times$$32$(Chrabaszcz et al., 2017)), and freeze all lower and middle layers (before Layer4.0.conv1 in ResNet-18). For regulating the feature maps, we select vital feature maps with a ratio $R=15$% in each layer after Layer4.0.conv1 (included). Hyperparameter $\beta$ in Equation 4 equals $0.0002$. Although our method doesn’t need to store samples in storage, for a fair comparison, we set the same buffer size to compare LightCL  with replay methods.
Experiment Details: We operate our experiments on a server equipped with an NVIDIA A800 GPU in Section 6.2 and Section 6.3. We further verify our method’s effectiveness on an edge device, Jetson Nano, in Section 6.4. All methods employ ResNet-18(He et al., 2016) with a learning rate of $0.01$. We set other hyperparameters as suggested in (Buzzega et al., 2020; Wang et al., 2022) for other CL  methods. For a fair comparison, all other methods employ the same pre-trained model in Section 6.2.1. Although the pre-trained model is more practical for application, most of CL  methods don’t take this into account and train the model from scratch. Hence, we test all methods without the pre-trained model to evaluate the effectiveness of LightCL  in Section 6.2.2. In order to compare with SparCL, we prune our model with sparsity ratio $90\%$ as the same as SparCL. Specifically, LightCL  uses static filter pruning for ResNet-18 and sets sparsity ratio as follows: 1) before Layer3.0 as $20\%$; 2) Layer3.0 and Layer3.1 as $70\%$; 3) others as $80\%$. Sparse is not the key point in our paper, and we set it up this way because we want to retain enough knowledge in the frozen layers.
6.2. Main Result
6.2.1. Comparison with CL  methods on the pre-trained model.
We evaluate results on Split CIFAR-10 and Tiny-ImageNet with a pre-trained ResNet-18 model under both CIL and TIL settings. The compared methods are divided into normal CL  methods (EWC, PNN, FDR, ER, DER++) and an efficient CL  method (SparCL). As shown in Table 1, the accuracy of LightCL  without sparsity increases compared with ER, which performs best in normal CL  methods except PNN. On Split Tiny-ImageNet, FLOPs of LightCL  greatly reduce at most $3.07$$\times$ and Memory footprint at most $4.79$$\times$, compared with normal CL  methods. That’s because LightCL  reduces the computation of backward process, and memory footprint of gradient and intermediate feature maps in frozen layers. Although the architecture-based method PNN has excellent accuracy under the TIL setting, it has tremendous FLOPs and memory footprint usage and also ignores the real application always under the CIL setting where the task’s identity is not known during testing. On the other hand, compared with the efficient method SparCL, the delaying performance of LightCL  with the same sparsity ratio is still competitive. Note that we employ a simple sparsity strategy without any optimization, such as dynamic sparsity(Wang et al., 2022). Surprisingly, the memory footprint is reduced at most $6.19$$\times$ compared with SparCLDER++. Although the FLOPs of LightCL  is higher than SparCL, LightCL  is faster than SparCL in the real application since SparCL still costs abundant time on memory transfer. This will be discussed in Section 6.4.
6.2.2. Comparison from scratch.
Although deploying a pre-trained model on edge devices is more practical, most CL  methods ignore it and focus more on CL  from scratch. To better illustrate the effectiveness of LightCL, we also compare LightCL  with other CL  methods in the setting of CL  from scratch. Since LightCL  needs to freeze the lower and middle layers, we take the first task in the dataset as a pre-trained process, and we freeze the lower and middle layers in the following tasks. From Table 2, LightCL  is still outstanding compared with other methods. LightCL0.9 with sparsity ratio $90\%$ also shows better performance compared with the SparCL. Although replay methods (FDR, ER, DER++) are ineffective with a limited buffer size of $15$, LightCL  is favorable in a limited buffer size.
6.3. Ablation
In this section, we illustrate the effectiveness of Maintaining Generalizability and Memorize Feature Patterns of LightCL  in Figure  4.
Discussion of Maintaining Generalizability: Maintaining Generalizability is an easy but effective method to delay forgetting and reduce resource usage. From Figure  4, compared with red and blue bars, all CL  methods can enhance performance with the pre-trained model. Most of them (SGD, EWC, FDR, ER, LightCL) can even further delay the forgetting problem by Maintaining Generalizability. Note that the accuracy of SparCL decreases when using Maintaining Generalizability. This phenomenon is because frozen layers weaken the dynamic pruning process, leading to the accuracy drop. Although most of CL  methods can benefit from Maintaining Generalizability, LightCL  still performs better than other methods.
Figure 4. Ablation experiments on Split CIFAR10 dataset under the TIL setting across 5 runs. Red bars denote the setting without the pre-trained model and Maintaining Generalizability (freezing layers). Blue bars denote the setting only with the pre-trained model. Green bars have both of them.
Discussion of Memorize Feature Patterns: We compare the performance of LightCL  with and without Memorize Feature Patterns. As shown in Figure  4, the green bars of SGD and LightCL  are the LightCL  with and without Memorize Feature Patterns, respectively. We can see that LightCL  with Memorize Feature Patterns can further increase the accuracy, proving its effectiveness. Let’s consider the resource consumption of this method. We need to store the Feature Standard $FS$, mentioned in Section 5.2, as a reference to regulate. Since we only need to store important feature maps after Layer4.0.conv1 (included), the overall consumption of memory footprint, suppose all feature maps need to be stored, occupies about $1.9$MB (on Split CIFAR-10), which is pretty small compared to what we reduce by Maintaining Generalizability.
6.4. Verification on Edge Device
Although we theoretically calculate the peak memory footprint for model parameters, gradients, and intermediate feature maps, some real-applied consumption, such as extra consumption in loading datasets, memory fragmentation, and some additional intermediate variables, also need to be considered. Hence, we deploy LightCL  on an edge device, Jetson Nano, to verify the efficiency in the real environment. We compare LightCL  with SparCLDER++. As shown in Figure  5, although the peak memory of them is all greater than the theoretical value shown in Table 1, the peak memory of LightCL is still greatly smaller compared to SparCL. Note that the consumption of the memory footprint is not always equal to the peak memory footprint. We also need to observe the reserved memory, red line in Figure  5. Since the peak memory of LightCL  is smaller than SparCL, the reserved memory is also smaller than SparCL. From the x-axis that represents the change in time, the training process of LightCL  is faster than SparCL since the memory transfer consumption has greatly hidden the training speed in SparCL. All of these findings suggest the effectiveness of LightCL in edge devices.
Figure 5. We deploy LightCL (top figure) and SparCLDER++ (bottom figure) on Jetson Nano and evaluate the dynamic change of memory footprint during CL. Only $5$ iterations are selected.
7. Conclusion
This paper proposes an efficient CL  algorithm LightCL  to overcome the catastrophic forgetting problem in edge scenarios. It is the first study to explore the distribution of generalizability in neural networks by evaluating two designed metrics of learning plasticity and memory stability. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits significant variation. Thus, we proposed LightCL  with Maintaining Generalizability and Memorize Feature Patterns. Experiments demonstrate that LightCL  is outstanding beyond other state-of-the-art   CL  methods with fewer resources, especially a lower memory footprint. Also, our method performs more favorably on edge devices and shows significant potential for efficient CL.